OpenStudy (anonymous):

Livestock feed. Soybean meal is 16% protein and corn meal is 9% protein. How many pounds of each should be mixed to get a 350-lb mixture that is 12% protein?

OpenStudy (saifoo.khan):

soybean: 150 lbs cornmeal: 200 lbs

OpenStudy (anonymous):

S+C=350 0.16*S+0.09*C=0.12*350=42 C=350-S 0.16S+(0.09(350-S))=42 0.07S+31.5=42 0.07S=10.5 S=10.5/0.07=150lbs C=350-150=200lbs

OpenStudy (anonymous):

ok thats the answer i came up with i wasn't sure thank you